given problem m = 3/4 , (1,-2)
y-(-2) = 3/4(x-1)
i don't know how to correctly solve this

Answer by mananth(16946)

I presume you need equation of line
..
m=3/4 passing through point (1,-2)
..
y=mx+b
first we find b by plugging the values of x,y,m in the equation
-2=3/4*1 +b
-2-3/4 =b
-11/4= b
...
y=3x/4 -11/4 is the equation of the line passing through (1,-2)
...
m.ananth@hotmail.ca